(x+3)(2x-1)(x-0.5)=0
Answer by stanbon(75887)

(x+3)(2x-1)(x-0.5)=0
---
Since the product is zero, one of the
factors must be zero.
---
x = -3 or x = 1/2 of x = 0.5
=================
cheers,
Stan H.